OpenStudy (anonymous):

According to the rule of 72, if the GDP of the Apex Federation is growing at 1.3% per year, its economy will double in approximately how many years? A. 55 years B. 17 years C. 42 years D. 41 years

OpenStudy (anonymous):

55 years

OpenStudy (anonymous):

72/1.3=55.38 so 55

OpenStudy (anonymous):

55 Years

OpenStudy (anonymous):

so wat did ya get on da test

OpenStudy (perl):

yep

OpenStudy (wolf1728):

If you want to know when its economy will actually double, use this formula: Years = {log(total) -log(Principal)} / (log(1 + rate)) we'll say total =2 and principal =1 (for the doubling) Years = {log(2) -log(1)} / (log(1 + .013)) Years = {log(2)} / (log(1.013)) Years = {log(2)} / (log(1.013)) Years = 0.3010299957 / 0.0056094454 Years = 53.6648414112 Years Wow it's surprisingly close to the rule of 72 calculation.

OpenStudy (wolf1728):

So I guess you agree Perl?

OpenStudy (anonymous):

so what now

OpenStudy (perl):

now we can ... i dont know

OpenStudy (perl):

@wolf1728 where did you get that formula, yes i agree

OpenStudy (wolf1728):

I got the formula from my website: http://www.1728.org/compint2.htm It has the interest formula solved for all 4 variables.

OpenStudy (anonymous):

thanks guys @perl @wolf1728 @tyga24 @kaylamarie081

OpenStudy (perl):

thanks wolf

OpenStudy (wolf1728):

u r welcome ar!3

OpenStudy (perl):

yw

OpenStudy (anonymous):

ur welcome